madz2345:

What is the sum of 10 and 15 divided by 2?

Lura:

here the sw: 10+15=25/2=12.5

Lura:

yw ^^

madz2345:

@lura wrote:
here the sw: 10+15=25/2=12.5
THANK YOU!! THIS IS GONNA HELP ON MY NEXT FEW QUESTIONS!!!

Lura:

:DDDDDD yw read dms lol

Lura:

scroll up in dms-

madz2345:

@lura wrote:
scroll up in dms-
XDD